# RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes

### v1

@@ 5.2 Implementation Details
5.3 Quantitative Assessment Results
@@
Results for Semantic Change Detection on RB-SCD Dataset: Table  4 presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics. As shown in the table, we have defined 11 types of semantic changes along with one background category, aiming to comprehensively cover the diverse real-world scenarios of road and bridge changes. These change types include both relatively straightforward categories—such as water to bridge and farmland to road, which typically involve simple backgrounds or abundant training samples—and more challenging categories, such as road to vegetation and grass to road, which are characterized by complex backgrounds, ambiguous boundaries, and limited sample sizes. Across this diverse range of semantic changes, our method achieves significant advantages in both the IoU and F1 metrics. Specifically, for the easily detectable categories—water to bridge, bridge to water, water to road, and farmland to road—our model attains IoU scores of 86.83%, 83.58%, 71.68%, and 68.40%, respectively, significantly outperforming other comparison methods. These results highlight the strong capability of our model in semantic discrimination and spatial reasoning. More importantly, our method also performs excellently on several challenging categories, including road to vegetation, road to grass, road to bare land, and grass to road, which suffer from sample scarcity and semantic ambiguity. These types of changes are often affected by similar surface textures, seasonal variability, and complex contextual information. Traditional methods are prone to false positives and missed detections under such conditions, whereas our model consistently demonstrates strong robustness and generalization ability by accurately identifying these subtle semantic transformations.
@@
Furthermore, we evaluate the performance of the latest RS foundation model, LWGANet (Lu et al., 2025), on our proposed dataset (as shown in Table 5). The results further underscore the challenging nature of our dataset. By incorporating fine-grained and practically meaningful semantic change categories in traffic scenarios, our dataset provides a more rigorous and instructive benchmark for training and evaluating CD models in real-world applications.
@@
Results for Binary Change Detection on RB-SCD Dataset: As shown in Table LABEL:tab:BCD, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics. In terms of the IoU metric, our model ranks first with a score of 65.83%, slightly outperforming ELGC-Net (Noman et al., 2024), which achieves 65.53%. This indicates that MFDCD demonstrates the best performance in terms of overlap between the predicted and ground truth change regions, enabling more precise localization of change areas. Regarding the F1, our MFDCD model also leads with a score of 79.39%, followed closely by ELGC-Net at 79.18%. Since the F1 is a harmonic mean of Pre and Rec, this result suggests that MFDCD achieves an excellent balance between detection accuracy and completeness. For recall, USSFC-Net (Lei et al., 2023) achieves the highest score of 75.07% among all models, while our model follows closely with a recall of 74.13%, ranking second. This demonstrates that USSFC-Net has a strong ability to detect all change areas, and our model also exhibits competitive performance in capturing changed regions. In terms of precision, our model outperforms others with a score of 85.47%, while ELGC-Net follows at 85.15%. This implies that our model has a lower false positive rate when identifying change areas. Overall, the results of the BCD task further indicate that existing methods still have room for improvement when applied to our RB-SCD dataset.
@@
Results on Public CD Datasets: To further prove the validity of our model, we also compare our proposed MFDCD with thirteen state-of-the-art CD techniques. To ensure a fair comparison, we reproduce these methods using publicly available code with default parameters. Table 7 presents the quantitative evaluations of various methods on three public CD datasets, including IoU, F1, Rec, and Pre metrics. As observed in Table 7, our method achieves outstanding results across all three publicly available datasets. Specifically, on the WHU-CD dataset, our approach outperforms the second-best method by 2.69% and 1.5% in IoU and F1, respectively. On the SYSU-CD dataset, it achieves improvements of 4.43% and 3.05%, while on the LEVIR-CD dataset, it surpasses the second-best method by 1.06% and 0.62%. These results further demonstrate the effectiveness and robustness of our method.
@@
As shown in Figure 7, we compared the visualized results of the proposed MFDCD with other state-of-the-art methods on RB-SCD dataset. From the figure, it can be observed that our method outperforms others in identifying small semantic change regions. For instance, as shown in the third row, our method better distinguishes between changes from grass to road and from vegetation to road—two types of changes with similar characteristics. Regarding pseudo-changes caused by background composition, such as in the fifth row, where a boat passing under a bridge with its bow color similar to the bridge’s edge color, both DEFO (Li et al., 2024) and ChangeCLIP (Dong et al., 2024) made incorrect predictions, while our method accurately detected the changes. Additionally, the proposed dataset presents significant challenges. For example, in the first row, all methods struggle to detect semantic changes under road shadows. As shown in the second row, farmland in autumn and winter appears visually similar to bare land, and in the fourth row, farmland in spring and summer shares similar colors with water, resulting in low inter-class differences. In such challenging scenes, other methods show many false detections, while our method performs slightly better but still leaves room for improvement.
@@
When comparing the BCD results of each method with the ground truth in Figure  8, it can be observed that for the unchanged areas (black regions representing true negatives), most models show relatively accurate results, with fewer false positives. In contrast, for the changed areas, the performance varies among models. Some models exhibit noticeable missed detections (blue regions representing false negatives), indicating they fail to identify all actual changes. Many models are capable of detecting building—related changes, which are common in the dataset. For example, in areas where buildings are newly constructed or modified, most models can capture at least part of the change, though the precision of boundary detection varies. In regions where there is a transition from bare land to road, our MFDCD and some others show better performance in accurately delineating the change area. Some methods, however, have varying degrees of missed detections. For instance, methods such as BIT (Chen et al., 2022) and ChangeFormer (Bandara and Patel, 2022) seem to have less accurate delineation compared to our method and ELGC-Net (Noman et al., 2024). In areas where buildings are under construction in one image and completed in another, different models have different performances. Some models, like ICIF (Feng et al., 2022) and TFI-GR (Li et al., 2022), can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives). Other models may miss parts of the change. In areas where the change from bare land to road is affected by shadows (resulting in pseudo-changes), most models struggle. Except for our MFDCD, which shows relatively better performance in distinguishing true changes from pseudo-changes caused by shadows, other models either misclassify the shadow areas as changes (false positives in red) or miss the actual changes (false negatives in blue).
@@
Table 8: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets. All values are reported as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
Table 9: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
Table 10: Effectiveness of graph filters with different numbers $k$ in TFF. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
5.5 Ablation Study
@@
To demonstrate the effectiveness of our proposed modules, we conduct extensive experiments on RB-SCD and three public datasets. As shown in Table 8, when the DFC module and TFF module are added to the baseline model, the mIoU on RB-SCD dataset increased by 5.44% and 4.78%, respectively. Additionally, the IoU on the LEVIR-CD, SYSU-CD, and WHU-CD datasets improved by 1.6%/2.12%/2.63% and 1.32%/1.91%/1.85%, respectively. As shown in Table 9, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module. Initially, we experiment with different coupling methods for ordinary features and frequency domain features. By comparing rows (b) and (c) with row (h), we can observe that ASFF and BTFF effectively dynamically couple visual features with frequency domain features, achieving the best results. Comparing rows (d) and (f), it can be seen that combining the low-level features at time $T_{1}$ with the corresponding high-frequency features at time $T_{2}$, due to semantic mismatches, results in worse performance compared to our designed ASFF. Comparing rows (e) and (g), it is clear that our designed BTFF captures the global structural changes across time, outperforming the coupling of high-level features and low-frequency features at the same time point. As shown in Table 10, we conduct ablation experiments on the number of layers in the frequency filter of TFF. As the value of $k$ increased, the model’s performance improved. However, when the number of filter layers reached six, the performance stabilized or slightly declined. Subsequent experiments are based on the optimal setting, $k=5$. To further demonstrate the effectiveness of our modules, we visualized the intermediate layer features of the network as heatmaps. As shown in Figure 9, the proposed modules effectively enhance the detection capability for roads and bridges.
6 Conclusion
@@
In this work, we introduce RB-SCD dataset, a SCD dataset specifically designed for complex traffic scenarios. The dataset features fine-grained pixel-level annotations across 11 semantic change categories and one background class, encompassing a wide range of road and bridge types. It also covers diverse geographic regions across multiple countries and urban environments, making it become a comprehensive and challenging benchmarks to date for road and bridge change analysis in RS imagery. Notably, RB-SCD supports both BCD and SCD tasks, offering flexibility for different application scenarios. By incorporating both common and rare semantic transformations, RB-SCD provides a valuable resource for the development and evaluation of CD models in real-world, high-density traffic contexts.
@@
In future work, we plan to extend our exploration of road and bridge change detection by integrating additional modalities such as elevation data or temporal sequences, and further investigating the role of frequency-domain representations in improving the interpretability and adaptability of CD models. As the label imbalance analysis and challenges for RB-SCD dataset discussed in Section  3.2, we also plan to expand the scale of the dataset and enrich it with more detailed spatiotemporal annotations to support broader research tasks such as urban infrastructure evolution and intelligent transportation systems. Furthermore, while MFDCD achieves strong performance, it is not yet lightweight enough for deployment in resource-constrained environments. Future research will focus on model compression and efficient architecture design to improve its applicability in real-time and edge scenarios.
CRediT authorship contribution statement
Qingling Shu: Writing—review & editing, Writing—original draft, Methodology, Funding acquisition, Formal analysis, Data curation. Sibao Chen: Software, Resources, Project administration. Zhihui You: Validation, Supervision. Wei Lu: Validation, Supervision. Jin Tang: Validation, Supervision. Bin Luo: Validation, Supervision.
Declaration of competing interest
The authors declare that they have no known competing finan- cial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgement
This work was supported in part by NSFC Key Project of International (Regional) Cooperation and Exchanges (No. 61860206004), NSFC Key Project of Joint Fund for Enterprise Innovation and Development (No. U20B2068, U24A20342) and National Natural Science Foundation of China (No. 61976004).

### v2

@@ V-B Implementation Details
V-C Quantitative Assessment Results
@@
Results for Semantic Change Detection on RB-SCD Dataset: Tab. IV presents the comparative results of our method and several advanced SCD methods across multiple evaluation metrics. As shown in the table, we have defined 11 types of semantic changes along with one background category, aiming to comprehensively cover the diverse real-world scenarios of road and bridge changes. These change types include both relatively straightforward categories—such as water to bridge and farmland to road, which typically involve simple backgrounds or abundant training samples—and more challenging categories, such as road to vegetation and grass to road, which are characterized by complex backgrounds, ambiguous boundaries, and limited sample sizes. Across this diverse range of semantic changes, our method achieves significant advantages in both the IoU and F1 metrics. Specifically, for the easily detectable categories—water to bridge, bridge to water, water to road, and farmland to road—our model attains IoU scores of 86.83%, 83.58%, 71.68%, and 68.40%, respectively, significantly outperforming other comparison methods. These results highlight the strong capability of our model in semantic discrimination and spatial reasoning. More importantly, our method also performs excellently on several challenging categories, including road to vegetation, road to grass, road to bare land, and grass to road, which suffer from sample scarcity and semantic ambiguity. These types of changes are often affected by similar surface textures, seasonal variability, and complex contextual information. Traditional methods are prone to false positives and missed detections under such conditions, whereas our model consistently demonstrates strong robustness and generalization ability by accurately identifying these subtle semantic transformations.
@@
Furthermore, we evaluate the performance of the latest RS foundation model, LWGANet [47], on our proposed dataset (as shown in Tab. V). The results further underscore the challenging nature of our dataset. By incorporating fine-grained and practically meaningful semantic change categories in traffic scenarios, our dataset provides a more rigorous and instructive benchmark for training and evaluating CD models in real-world applications.
@@
Results for Binary Change Detection on RB-SCD Dataset: As shown in Tab. VI, in the BCD task based on the RB-SCD dataset, different models exhibit varying performance across different evaluation metrics. In terms of the IoU metric, our model ranks first with a score of 65.83%, slightly outperforming ELGC-Net [50], which achieves 65.53%. This indicates that MFDCD demonstrates the best performance in terms of overlap between the predicted and ground truth change regions, enabling more precise localization of change areas. Regarding the F1, our MFDCD model also leads with a score of 79.39%, followed closely by ELGC-Net at 79.18%. Since the F1 is a harmonic mean of Pre and Rec, this result suggests that MFDCD achieves an excellent balance between detection accuracy and completeness. For recall, USSFC-Net [49] achieves the highest score of 75.07% among all models, while our model follows closely with a recall of 74.13%, ranking second. This demonstrates that USSFC-Net has a strong ability to detect all change areas, and our model also exhibits competitive performance in capturing changed regions. In terms of precision, our model outperforms others with a score of 85.47%, while ELGC-Net follows at 85.15%. This implies that our model has a lower false positive rate when identifying change areas. Overall, the results of the BCD task further indicate that existing methods still have room for improvement when applied to our RB-SCD dataset.
@@
Results on Public CD Datasets: To further prove the validity of our model, we also compare our proposed MFDCD with thirteen state-of-the-art CD techniques. To ensure a fair comparison, we reproduce these methods using publicly available code with default parameters. Tab. VII presents the quantitative evaluations of various methods on three public CD datasets, including IoU, F1, Rec, and Pre metrics. As observed in Tab. VII, our method achieves outstanding results across all three publicly available datasets. Specifically, on the WHU-CD dataset, our approach outperforms the second-best method by 2.69% and 1.5% in IoU and F1, respectively. On the SYSU-CD dataset, it achieves improvements of 4.43% and 3.05%, while on the LEVIR-CD dataset, it surpasses the second-best method by 1.06% and 0.62%. These results further demonstrate the effectiveness and robustness of our method.
@@
As shown in Fig. 7, we compared the visualized results of the proposed MFDCD with other state-of-the-art methods on RB-SCD dataset. From the figure, it can be observed that our method outperforms others in identifying small semantic change regions. For instance, as shown in the third row, our method better distinguishes between changes from grass to road and from vegetation to road—two types of changes with similar characteristics. Regarding pseudo-changes caused by background composition, such as in the fifth row, where a boat passing under a bridge with its bow color similar to the bridge’s edge color, both DEFO [35] and ChangeCLIP [23] made incorrect predictions, while our method accurately detected the changes. Additionally, the proposed dataset presents significant challenges. For example, in the first row, all methods struggle to detect semantic changes under road shadows. As shown in the second row, farmland in autumn and winter appears visually similar to bare land, and in the fourth row, farmland in spring and summer shares similar colors with water, resulting in low inter-class differences. In such challenging scenes, other methods show many false detections, while our method performs slightly better but still leaves room for improvement.
@@
When comparing the BCD results of each method with the ground truth in Fig. 8, it can be observed that for the unchanged areas (black regions representing true negatives), most models show relatively accurate results, with fewer false positives. In contrast, for the changed areas, the performance varies among models. Some models exhibit noticeable missed detections (blue regions representing false negatives), indicating they fail to identify all actual changes. Many models are capable of detecting building—related changes, which are common in the dataset. For example, in areas where buildings are newly constructed or modified, most models can capture at least part of the change, though the precision of boundary detection varies. In regions where there is a transition from bare land to road, our MFDCD and some others show better performance in accurately delineating the change area. Some methods, however, have varying degrees of missed detections. For instance, methods such as BIT [18] and ChangeFormer [19] seem to have less accurate delineation compared to our method and ELGC-Net [50]. In areas where buildings are under construction in one image and completed in another, different models have different performances. Some models, like ICIF [17] and TFI-GR [48], can detect the change but with somewhat blurred boundaries compared to the ground truth (white regions representing true positives). Other models may miss parts of the change. In areas where the change from bare land to road is affected by shadows (resulting in pseudo-changes), most models struggle. Except for our MFDCD, which shows relatively better performance in distinguishing true changes from pseudo-changes caused by shadows, other models either misclassify the shadow areas as changes (false positives in red) or miss the actual changes (false negatives in blue).
@@
TABLE VIII: Ablation studies of DFC and TFF when trained on RB-SCD and public CD datasets. All values are reported as percentages (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
TABLE IX: Effectiveness of the DFC module when trained on RB-SCD and public CD datasets. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
TABLE X: Effectiveness of graph filters with different numbers $k$ in TFF. All values are reported in percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
@@
TABLE XI: Effectiveness of MFDCD with Different Backbones when Trained on Four CD Datasets. All Values are Reported in Percentage (%). Results highlighted in red indicate the best performance, while those in blue denote the second-best performance.
Backbone Params (M) FLOPs (G) RB -SCD LEVIR -CD SYSU -CD WHU -CD
ResNet18 13.75 11.32 55.53 83.18 71.35 89.04
ResNet34 22.80 16.54 56.31 83.78 72.10 89.73
ResNet50 103.78 49.02 57.52 84.69 72.63 90.43
V-E Ablation Study
@@
To demonstrate the effectiveness of our proposed modules, we conduct extensive experiments on RB-SCD and three public datasets. As shown in Tab. VIII, when the DFC module and TFF module are added to the baseline model, the mIoU on RB-SCD dataset increased by 5.44% and 4.78%, respectively. Additionally, the IoU on the LEVIR-CD, SYSU-CD, and WHU-CD datasets improved by 1.6%/2.12%/2.63% and 1.32%/1.91%/1.85%, respectively. As shown in Tab. IX, we conduct detailed ablation studies on the ASFF and BTFF blocks in the DFC module. Initially, we experiment with different coupling methods for ordinary features and frequency domain features. By comparing rows (b) and (c) with row (h), we can observe that ASFF and BTFF effectively dynamically couple visual features with frequency domain features, achieving the best results. Comparing rows (d) and (f), it can be seen that combining the low-level features at time $T_{1}$ with the corresponding high-frequency features at time $T_{2}$, due to semantic mismatches, results in worse performance compared to our designed ASFF. Comparing rows (e) and (g), it is clear that our designed BTFF captures the global structural changes across time, outperforming the coupling of high-level features and low-frequency features at the same time point. As shown in Tab. X, we conduct ablation experiments on the number of layers in the frequency filter of TFF. As the value of $k$ increased, the model’s performance improved. However, when the number of filter layers reached six, the performance stabilized or slightly declined. Subsequent experiments are based on the optimal setting, $k=5$.
To investigate the impact of backbone networks on the performance of the MFDD model in the CD tasks, as shown in Tab.  XI, in terms of parameters and computational complexity, ResNet18 has the lowest cost (13.75M parameters, 11.32G FLOPs), while ResNet50 has the highest (103.78M parameters, 49.02G FLOPs). In terms of detection accuracy, for all four datasets, the deeper the backbone network, the higher the accuracy. However, a deeper network also leads to a doubling of parameters and computational complexity. In this paper, ResNet50 is selected as the backbone network.
To more clearly illustrate the influence of our proposed modules, we visualized heatmaps for these modules. As depicted in columns (b)/(h) and (c)/(i) of Fig. 9, the low-level and high-level features from the backbone network typically exhibit certain characteristics. The low-level features may contain fine-grained but scattered information, while the high-level features, though more abstract, might lose some detailed spatial information. In columns (d)/(j), (e)/(k), and (f)/(l), after undergoing the ASFF block, BTTF block and TFF module respectively, the feature maps show progressive changes. The ASFF block helps to fuse features in a reasonable way, making the information more integrated. The BTTF block further refines the features, enhancing the relevant change information. The TFF module finally outputs feature maps that can better highlight the change regions, eliminating some irrelevant noise and enabling the acquisition of more distinct change-related information, which showcases the efficacy of the introduced modules in feature processing for CD tasks.
VI Conclusion
@@
In this work, we introduce RB-SCD dataset, a SCD dataset specifically designed for complex traffic scenarios. The dataset features fine-grained pixel-level annotations across 11 semantic change categories and one background class, encompassing a wide range of road and bridge types. It also covers diverse geographic regions across multiple countries and urban environments, making it become a comprehensive and challenging benchmarks to date for road and bridge change analysis in RS imagery. Notably, RB-SCD supports both BCD and SCD tasks, offering flexibility for different application scenarios. By incorporating both common and rare semantic transformations, RB-SCD provides a valuable resource for the development and evaluation of CD models in traffic contexts.
@@
In future work, we plan to extend our exploration of road and bridge change detection by integrating additional modalities such as elevation data or temporal sequences, and further investigating the role of frequency-domain representations in improving the interpretability and adaptability of CD models. As the label imbalance analysis and challenges for RB-SCD dataset discussed in Sec. III-B, we also plan to expand the scale of the dataset and enrich it with more detailed spatiotemporal annotations to support broader research tasks such as urban infrastructure evolution and intelligent transportation systems. Furthermore, while MFDCD achieves strong performance, it is not yet lightweight enough for deployment in resource-constrained environments. Future research will focus on model compression and efficient architecture design to improve its applicability in real-time and edge scenarios.